OpenStudy (clamin):

Solve this equation by factoring::: 7x^2-21x-189= 7

OpenStudy (tmak11):

man why didn't i take algebra 2 in-staid of geometry ...if you had asked me this last year i could tell you in less than a minuet

OpenStudy (tmak11):

(x-? )(x-?) thats as far as i remember

TheSmartOne (thesmartone):

first take out 7 from it.

TheSmartOne (thesmartone):

Wait actually subtract 7 on both sides and then divide by 7.

TheSmartOne (thesmartone):

\[7x^2-21x-196=7(x^2-3x-28)=0\]

TheSmartOne (thesmartone):

Now try to factor it.

TheSmartOne (thesmartone):

@clamin

OpenStudy (clamin):

whats that in the parenthesis?

TheSmartOne (thesmartone):

I don't understand what you are asking...

OpenStudy (clamin):

how did you get 3x and 28??

TheSmartOne (thesmartone):

i took out 7 from the whole equation so that way it would be easier to factor

TheSmartOne (thesmartone):

your factors would be -7 and +4

TheSmartOne (thesmartone):

add -7 and +4 and you get -3 and multiply them you get -28

OpenStudy (clamin):

youre confusing

TheSmartOne (thesmartone):

Sorry